# An Effective Way to Determine the Separability of Quantum State

### v1

@@ Introduction
Figure 1: Entanglement detection criteria. (a) Horodecki’s $3\times 3$ bound entangled states with white noise $\rho_{\mathrm{H}}(s,p)$. (b) UPB bound entangled states with white noise $\rho_{\mathrm{UPB}}(p)$ and $50000$ randomly generated chessboard states $\rho_{\mathrm{CB}}$. Here, we simply take $t_{A}=t_{B}=t$. The larger $t$ detects the more entangled states as shown in (a) and (b), in which $t=\sqrt{3}/2$ and $t=\sqrt{3}$ correspond to CCNR and ESIC criteria, respectively. Diagrams indicate explicitly that the criterion given in Observation 1 outperform CCNR and ESIC criteria under the three typical classes of bound entangled states.
@@
Encouraged by the results in Fig.1, one may naturally ask whether larger $t$ means a more powerful criterion, or, specifically, whether the ESIC criterion is stronger than the CCNR or not, which was conjectured to be true in Ref. [5]. Unfortunately, Eq. 45 can not tell, since the left hand side (lhs) of it contains as well parameters $t_{A}$ and $t_{B}$. Nevertheless, we find this question can be readily addressed under the local filtering transformation (LFT), which preserves the separability of a given state, i.e. $\rho\xmapsto{}\tilde{\rho}=(F_{A}\otimes F_{B})\rho(F_{A}\otimes F_{B})^{\dagger}/\mathrm{Tr}[(F_{A}\otimes F_{B})\rho(F_{A}\otimes F_{B})^{\dagger}]$ with $F_{A/B}$ the arbitrary invertible matrices [51]. LFT transforms the full local rank state into a normal form with maximally mixed subsystems [51, 52]
@@
By dint of quantum states formulated in form of Eq. 25 the bipartite separability problem [52] will be solved. Concretely, we have the following theorem (see Supplemental Material for proof):
@@
Now that the lhs of Eq. 27 is independent of parameters $t_{A},t_{B}$, one can then strictly compare it with other criteria for different $t_{A}$ and $t_{B}$. Notice that the minimal value of $\kappa(t_{A},t_{B})$, i.e. $\min_{\{t_{A},t_{B}\}}\kappa(t_{A},t_{B})=2\sqrt{(d_{A}-1)(d_{B}-1)/d_{A}d_{B}}$, corresponds to the strongest constraint on separability, which reproduces the Proposition 6 of Ref. [21] when $d_{A}=d_{B}$. Furthermore, the CCNR and ESIC criteria will reexhibit respectively while we take $\kappa(t_{A},t_{B})=2-2/\sqrt{d_{A}d_{B}}$ and $4-2\sqrt{(d_{A}+1)(d_{B}+1)/d_{A}d_{B}}$, which also implies that in normal form the ESIC criterion is stronger than CCNR and they coincide as $d_{A}=d_{B}$.
Relations to entanglement witness
@@
An entanglement witness is a Hermitian operator which has nonnegative expectation values for all separable states and negative one for some entangled states [7, 53]. Entanglement witness provides an experimental friendly tool to analyse entanglement [3]. To this end, we reformulate the Theorem 1 in terms of entanglement witness. Given a correlation matrix $\mathcal{C}$ for measurement vectors $\boldsymbol{A},\boldsymbol{B}$, we readily find
$\|\mathcal{C}\|_{\mathrm{tr}}=\sum_{\mu=1}^{r}\braket{A^{\prime}_{\mu}\otimes B^{\prime}_{\mu}}\;.$ (29)
@@
Here, $A^{\prime}_{\mu}=\sum_{\nu}P_{\nu\mu}A_{\nu}$, $B^{\prime}_{\mu}=\sum_{\nu}Q_{\nu\mu}B_{\nu}$ and $P,Q$ achieve a singular value decomposition of $\mathcal{C}$, i.e. $\mathcal{C}=P\Sigma Q^{\mathrm{T}}$, where $P,Q$ are orthogonal matrices, $\Sigma=\operatorname{diag}\{\sigma_{1},\cdots,\sigma_{r},0,\cdots,0\}$ with singular values $\sigma_{k}$ and $\operatorname{rank}(\mathcal{C})=r$ [54]. Thus, one may immediately get an inequality of correlation function similar to CHSH inequality Eq. 6 for separable state
@@
This means that based upon the extended Bloch space, entanglement witness will be simply generated for generic observables.
Further extensions
Through the construction of a measurement-induced Bloch space, we have established a general condition for separability which arises from the constraint of the length of the MIBVs. However, not only the moduli of MIBVs but also their directions suffice a constraint. As an illustration, we consider the measurement vector $\boldsymbol{A}=\boldsymbol{B}=\vec{\pi}$, in which case the measurement-induced Bloch space reduces to the familiar Bloch space. Then the angel $\theta$ between any two Bloch vectors of pure states satisfies $-\frac{1}{d-1}\leq\cos\theta\leq 1$ [55, 35]. Given a $d\times d$ bipartite state, if it is separable, i.e. $\mathcal{T}=\sum_{i}p_{i}\vec{r}_{i}\vec{s}_{i}^{\mathrm{T}}$ with $\vec{r}_{i}$ and $\vec{s}_{i}$ being the Bloch vectors, we find the criterion
$-\frac{2}{d}\leq\mathrm{Tr}[\mathcal{T}]\leq\frac{2(d-1)}{d}$ (32)
exists, otherwise the entanglement survives. We notice $\mathrm{Tr}[\mathcal{T}]=\sum_{\mu=1}^{d^{2}-1}\braket{\pi_{\mu}\otimes\pi_{\mu}}$, so Eq. 32 yields the entanglement witness $\mathcal{W}=\frac{2}{d}\mathds{1}\otimes\mathds{1}+\sum_{\mu}\pi_{\mu}\otimes\pi_{\mu}$. Considering arbitrary dimensional Werner states [47, 56]
$\rho_{\mathrm{W}}=\frac{1}{d^{2}}\mathds{1}\otimes\mathds{1}+\frac{1}{4}\sum_{\mu=1}^{d^{2}-1}\frac{2(d\phi-1)}{d(d^{2}-1)}\pi_{\mu}\otimes\pi_{\mu}\;,\;\forall\phi\in[-1,1]\;,$
which is separable if and only if $\phi\geq 0$ [47]. An simple calculation shows that Eq. 32 is a necessary and sufficient condition taking effect in full parameter space $\phi\in[-1,0)$ , while CCNR and ESIC criteria are valid merely in $\phi\in[-1,-\frac{d-2}{d}]$. With the growth of dimensionality, the entanglement detectable intervals of CCNR and ESIC criteria tend to none. This example tells that we can establish more powerful separability condition than existing ones by only employing the inherent constraints in Bloch space. On this account, we are tempting to conclude that the inherent constraints in Bloch space may address a lot of the separability issues.
Concluding remarks
@@ Acknowledgements
This work was supported in part by the National Natural Science Foundation of China(NSFC) under the Grants 11975236, 12235008, the Fundamental Research Funds for the Central Universities and by the University of Chinese Academy of Sciences.
Supplemental Material
Here we elaborate and derive in detail the main results discussed in the main text.
The proof of Theorem 1
Given arbitrary measurement vectors $\boldsymbol{A}$ and $\boldsymbol{B}$ with with $\sum_{\mu}t_{\mu}^{A}\vec{a}_{\mu}=\sum_{\mu}t_{\mu}^{B}\vec{b}_{\mu}=0$, the correlation matrices of the separable states can be expressed as
@@
Here, $\vec{\alpha}_{i}$ and $\vec{\beta}_{i}$ are the generalized Bloch vectors associated with $\boldsymbol{A}$ and $\boldsymbol{B}$ as defined in the main text. Employing the convexity property of trace norm $\|X\|_{\operatorname{tr}}=\mathrm{Tr}\left[\sqrt{X^{\dagger}X}\right]$, we have
@@
Here, we make use of $\sum_{\mu}t_{\mu}^{A}\vec{a}_{\mu}=\sum_{\mu}t_{\mu}^{B}\vec{b}_{\mu}=0$ and ||$\cdot$|| ⋅ | denotes Euclidean norm of a real vector. So, the correlation matrices of the separable states are limited by the size of the measurement-induced Bloch space (MIBS). Next, we prove that the MIBS is bounded and its boundary can be obtained simply.
Given a measurement vector $\boldsymbol{A}$, quantum states can be transformed into the measurement-induced Bloch vector (MIBVs) which constitute the MIBS. It is found that the MIBVs satisfy the following equation [1]
$\vec{\alpha}^{\mathrm{T}}\Omega^{-}\vec{\alpha}=\mathrm{Tr}[\rho^{2}]-\frac{1}{d}\;.$ (36)
Here, $\Omega_{\mu\nu}\equiv\mathrm{Tr}[\bar{A}_{\mu}\bar{A}_{\nu}]=2\vec{a}_{\mu}\cdot\vec{a}_{\nu}$ and $\Omega^{-}$ denotes the Moore-Penrose inverse of $\Omega$. Eq. 36 is an ellipsoid equation, which characterizes the MIBS, that is,
$\mathcal{B}(\boldsymbol{A})=\left\{\vec{\alpha}\Big{|}\sqrt{\frac{d}{d\mathrm{Tr}[\rho^{2}]-1}}\vec{\alpha}^{\mathrm{T}}\Omega^{-}\sqrt{\frac{d}{d\mathrm{Tr}[\rho^{2}]-1}}\vec{\alpha}\leq 1\right\}\;.$ (37)
Matrix $\Omega$ completely depicts the properties of the ellipsoid defined by Eq. 37, such as the lengths of the semi-axes are given by $\sqrt{\frac{(d\mathrm{Tr}[\rho^{2}]-1)\lambda_{i}}{d}}$, with $\lambda_{i}$ being the eigenvalues of $\Omega$; the eigenvectors of $\Omega$ determines the directions of the semi-axes [2]. Obviously, Bloch space is a special case of $\boldsymbol{A}=\vec{\pi}$. So, for arbitrary measurement vector $\boldsymbol{A}$, we have
$|\vec{\alpha}|^{2}\leq\frac{d-1}{d}\lambda_{\max}(\Omega)\;,$ (38)
where $\lambda_{\max}(X)$ is the maximal eigenvalue of matrix $X$. Finally, we prove
$\|\mathcal{C}\|_{\operatorname{tr}}\leq\kappa_{A}\kappa_{B}\;.$ (39)
Here, $\kappa_{A}=\sqrt{(\vec{t}\,^{A})^{2}/d_{A}^{2}+\frac{d_{A}-1}{d_{A}}\lambda_{\max}(\Omega_{A})}$ and $(\Omega_{A})_{\mu\nu}=2\vec{a}_{\mu}\cdot\vec{a}_{\nu}$, similarly for $\kappa_{B}$.
The proof of Observation 1

### v2

@@ 4 The separability criteria
Figure 1: Entanglement detection for three classes of bound entangled states. (a) Horodecki’s $3\times 3$ bound entangled states with white noise $\rho_{\mathrm{H}}(s,p)$. (b) UPB bound entangled states with white noise $\rho_{\mathrm{UPB}}(p)$ and $50000$ randomly generated chessboard states $\rho_{\mathrm{CB}}$. The larger $t$ detects the more entangled states as shown in (a) and (b), in which $t=\sqrt{3}/2$ and $t=\sqrt{3}$ correspond to CCNR and ESIC criteria, respectively. Diagrams indicate explicitly that the criterion given in Observation 1 outperform CCNR and ESIC criteria under the three typical classes of bound entangled states.
In Fig. 1, we compare the performance of 1, the CCNR criterion, and the ESIC criterion for three classes of bound entangled states. Here, a larger value of $t$ indicates greater robustness against white noise, enabling the detection of a larger portion of entangled states. Fig. 1(a) presents results for Horodecki’s bound entangled states with white noise, while Fig. 1(b) details the detection results for the UPB state with white noise (solid peru line) and $50000$ randomly generated chessboard states (blue dashed line). Notably, as illustrated in the figure, 1 reproduces the CCNR and ESIC criteria as two special cases when $t=\sqrt{3}/2$ and $t=\sqrt{3}$, respectively.
When the two involved subsystems share the same dimension, an enhanced separability condition is readily established (see Appendix D for the proof):
Observation 2.
Given $\boldsymbol{A}=\boldsymbol{B}$, let their trace parts satisfy $t_{\mu}^{A}=t_{\mu}^{B}=t$ and traceless parts correspond to a regular simplex. The correlation matrix of the $d\times d$ separable state $\rho$ then satisfies
$\|\mathcal{C}\|_{\mathrm{tr}}\leq t^{2}+\frac{2d}{d+1}\;,$ (18)
and
$\mathrm{Tr}[\mathcal{C}]\geq t^{2}-\frac{2d}{d^{2}-1}\;,$ (19)
otherwise, $\rho$ is entangled. Here, $t$ is arbitrary real parameter.
We now compare the criterion proposed by Sarbicki et al. [19] with 2. To this end, we first consider the random $2\times 2$, $3\times 3$ and $5\times 5$ states with the Hilbert-Schmidt distribution [^1] If a square random matrix $A$ whose entries are drawn independently according to the standard complex Gaussian distribution, then the density matrix $\rho=AA^{\dagger}/\mathrm{Tr}[AA^{\dagger}]$ obeys the Hilbert-Schmidt distribution [18].. In Fig. 2, we plot the fractions of entangled states identified by Eq. 18 (peru solid line) and the criterion proposed by Sarbicki et al. (blue solid line) for 50000 randomly generated $2\times 2$, $3\times 3$ and $5\times 5$ states following the Hilbert-Schmidt distribution. As the parameters $t$ and $h$ increase, the two inequalities tend to align, detecting the same fractions of entangled states. Therefore, we conjecture that the criterion proposed by Sarbicki et al. and 1 are equivalent when the entire parameter spaces are explored. Notably, a significantly larger percentage of entangled states is detected in the $3\times 3$ case compared to other dimensions, as highlighted in [18], underscoring the uniqueness of dimension $3$.
Figure 2: Entanglement detection for the randomly generated $2\times 2$, $3\times 3$ and $5\times 5$ states with the Hilbert-Schmidt distribution. With the parameters $t$ or $h$ increasing, the two inequalities tend to be agreeable and detect the same fractions.
Next, we specifically consider arbitrary dimensional Werner states [53, 60]
$\rho_{\mathrm{W}}\!=\!\frac{1}{d^{2}}\mathds{1}\otimes\mathds{1}\!+\!\frac{1}{4}\sum_{\mu=1}^{d^{2}-1}\frac{2(d\phi-1)}{d(d^{2}-1)}\pi_{\mu}\otimes\pi_{\mu}\;,$ (20)
where the parameter range is $\phi\in[-1,1]$. The Werner states are separable if and only if $\phi\geq 0$ [53]. For the arbitrary dimensional Werner states, Sarbicki et al.’s criteria give $|d\phi-1|\leq\sqrt{d-1+h_{A}^{2}}\sqrt{d-1+h_{B}^{2}}$, if $\rho_{\mathrm{W}}$ is separable. The optimal inequality constraint is achieved by $\min_{\{h_{A},h_{B}\}}\sqrt{d-1+h_{A}^{2}}\sqrt{d-1+h_{B}^{2}}=d-1$, which recognizes the entangled range $\phi\in[-1,-\frac{d-2}{d})$. Notably, the intervals of entanglement detectable by Sarbicki et al.’s criteria tend to diminish to zero as the dimensionality increases. A straightforward calculation demonstrates that Eq. 19 serves as a necessary and sufficient condition applicable across the entire parameter space $\phi\in[-1,0)$ for the arbitrary dimensional Werner states.
5 Enhanced separability criteria and local filtering transformation
Encouraged by the results in Fig.1, one might naturally ask whether a larger $t$ indicates a more powerful criterion. Specifically, the question arises as to whether the ESIC criterion is stronger than the CCNR criterion, a conjecture proposed in Ref. [18]. Unfortunately, Eq. 17 cannot tell more, since the left hand side (LHS) of it contains as well parameters $t_{A}$ and $t_{B}$.
@@ 5 Enhanced separability criteria and local filtering transformation
Nevertheless, it is found that this question can be readily addressed under the local filtering transformation (LFT), which preserves the separability of a given state, i.e. $\rho\xmapsto{}\tilde{\rho}=(F_{A}\otimes F_{B})\rho(F_{A}\otimes F_{B})^{\dagger}/\mathrm{Tr}[(F_{A}\otimes F_{B})\rho(F_{A}\otimes F_{B})^{\dagger}]$ with $F_{A/B}$ the arbitrary invertible matrices [61]. LFT transforms the full local rank state into a normal form with maximally mixed subsystems [61, 62]
@@
It is proved that all $d_{A}\times d_{B}$ states with local ranks $n\leq d_{A}$ and $m\leq d_{B}$ are either reducible to $n\times m$ bipartite states with full local ranks, or entangled in Ref. [62]. Therefore we only need to focus on the separability problem of quantum states with full local ranks, which has been depicted in the flow diagram of Fig. 3.
Figure 3: The flow diagram of entanglement detection. The bipartite separability problem reduces to the bipartite states with full local ranks and further normal form states.
In this sense, the bipartite separability problem can be addressed through quantum states expressed in the form of Eq. 21. Specifically, we present the following theorem (see Appendix F for the proof):
@@
Now that the LHS of Eq. 23 is independent of parameters $t_{A},t_{B}$, we can strictly compare it with other criteria for different values $t_{A}$ and $t_{B}$. Notice that the minimal value of $\kappa(t_{A},t_{B})$, i.e. $\min_{\{t_{A},t_{B}\}}\kappa(t_{A},t_{B})=2\sqrt{(d_{A}-1)(d_{B}-1)/d_{A}d_{B}}$, corresponds to the strongest constraint on separability, which reproduces the Proposition 6 of Ref. [26] when $d_{A}=d_{B}$. After LFT, the CCNR and ESIC criteria correspond to $\kappa(t_{A},t_{B})=2-2/\sqrt{d_{A}d_{B}}$ and $4-2\sqrt{(d_{A}+1)(d_{B}+1)/d_{A}d_{B}}$ respectively. This indicates that in normal form, the ESIC criterion is stronger than CCNR, with both coinciding when $d_{A}=d_{B}$. In Table 2, we present a summary of how to obtain de Vicente’s criterion, Sarbicki et al.’s criterion, CCNR, ESIC, and Proposition 6 in [26] after applying local filtering transformations, highlighting the explicit relationships among them.
Table 2: Here, we summarize the relations between our criteria and all these mentioned in the main text after applying LFT. After applying LFT, for the optimal parameter values, 1 coincides with Sarbicki et al.’s criterion, i.e. $\|\widetilde{\mathcal{T}}\|_{\mathrm{tr}}\leq 2\sqrt{(d_{A}-1)(d_{B}-1)/d_{A}d_{B}}$, which reproduces the Proposition 6 of Ref. [26] when $d_{A}=d_{B}$ and de Vicente’s criterion. The ESIC criterion is stronger than CCNR and they coincide as $d_{A}=d_{B}$ after applying LFT.
Criteria $M_{A}/M_{B}$ $\mathcal{A}/\mathcal{B}$ $\kappa_{A}\kappa_{B}-\frac{|\vec{t}^{A}||\vec{t}^{B}|}{d_{A}d_{B}}$
de Vicente’s criterion $\left[\begin{smallmatrix}0&\vec{0}^{\mathrm{T}}\\ \vec{0}&\mathds{1}_{d^{2}-1}\end{smallmatrix}\right]$ $\mathds{1}_{d^{2}-1}$ $\sqrt{\frac{4(d_{A}-1)(d_{B}-1)}{d_{A}d_{B}}}$
Sarbicki et al.’s criterion $\left[\begin{smallmatrix}h&\vec{0}^{\mathrm{T}}\\ \vec{0}&\mathds{1}_{d^{2}-1}\end{smallmatrix}\right]$ $\mathds{1}_{d^{2}-1}$ $\sqrt{\frac{4(d_{A}-1+h_{A}^{2})(d_{B}-1+h_{B}^{2})}{d_{A}d_{B}}}-\frac{|h_{A}||h_{B}|}{d_{A}d_{B}}$
This work $\left[\begin{smallmatrix}\frac{t}{\sqrt{2d}}&\cdots&\frac{t}{\sqrt{2d}}\\ \vec{a}_{1}&\cdots&\vec{a}_{d^{2}}\end{smallmatrix}\right]$ $\frac{d^{2}}{d^{2}-1}\mathds{1}_{d^{2}-1}$ $\sqrt{t_{A}^{2}+\frac{2d_{A}}{d_{A}+1}}\sqrt{t_{B}^{2}+\frac{2d_{B}}{d_{B}+1}}-|t_{A}||t_{B}|$
CCNR $\left[\begin{smallmatrix}\frac{1}{\sqrt{d^{2}-1}}&\cdots&\frac{1}{\sqrt{d^{2}-1}}\\ \vec{a}_{1}&\cdots&\vec{a}_{d^{2}}\end{smallmatrix}\right]$ $\frac{d^{2}}{d^{2}-1}\mathds{1}_{d^{2}-1}$ $\frac{2(d_{A}d_{B}-\sqrt{d_{A}d_{B}})}{\sqrt{(d_{A}^{2}-1)(d_{B}^{2}-1)}}$
ESIC $\left[\begin{smallmatrix}\frac{1}{\sqrt{d-1}}&\cdots&\frac{1}{\sqrt{d-1}}\\ \vec{a}_{1}&\cdots&\vec{a}_{d^{2}}\end{smallmatrix}\right]$ $\frac{d^{2}}{d^{2}-1}\mathds{1}_{d^{2}-1}$ $\frac{4d_{A}d_{B}-2\sqrt{d_{A}d_{B}(d_{A}+1)(d_{B}+1)}}{\sqrt{(d_{A}^{2}-1)(d_{B}^{2}-1)}}$
Proposition 6 in [26] $\left[\begin{smallmatrix}0&\cdots&0\\ \vec{a}_{1}&\cdots&\vec{a}_{d^{2}}\end{smallmatrix}\right]$ $\frac{d^{2}}{d^{2}-1}\mathds{1}_{d^{2}-1}$ $\sqrt{\frac{4d_{A}d_{B}}{(d_{A}+1)(d_{B}+1)}}$
6 Relations to entanglement witness
@@
An entanglement witness is a Hermitian operator which has nonnegative expectation values for all separable states and negative one for some entangled states [7, 63]. Entanglement witness provides an experimental friendly tool to analyse entanglement [3]. We next reformulate the Theorem 1 in terms of entanglement witness. To this end, we provide a new understanding for the trace norm of correlation matrix for observables $m$-tuple $\boldsymbol{A}$ and $n$-tuple $\boldsymbol{B}$. Similar to the orbit of a set under group action [64], one can define the orbit of $m$-tuple $\boldsymbol{A}$ under the orthogonal group $O(m)$
$\mathcal{O}(\boldsymbol{A})=\{O\boldsymbol{A}|O\in O(m)\}\;.$ (25)
Here, $\boldsymbol{A}^{\prime}\in\mathcal{O}(\boldsymbol{A})$ implies $A^{\prime}_{\mu}=\sum_{\nu}O_{\mu\nu}A_{\nu},\mu=1,\cdots,m$. It is readily to prove that the all elements of $\mathcal{O}(\boldsymbol{A})$ yield the same MIBS up to a rotation. We can define the orbit $\mathcal{O}(\boldsymbol{B})$ of $n$-tuple $\boldsymbol{B}$ as well. Naturally, given $m$-tuple $\boldsymbol{A}$ and $n$-tuple $\boldsymbol{B}$, the orbit of the correlation matrix is defined as
$\mathcal{O}(\mathcal{C})\!\equiv\!\{\braket{\!A_{\mu}\otimes B_{\nu}\!}|\boldsymbol{A}\!\in\!\mathcal{O}(\boldsymbol{A}),\boldsymbol{B}\!\in\!\mathcal{O}(\boldsymbol{B})\!\}\;.$ (26)
Without loss of generality, assume $m\geq n$. By padding $\boldsymbol{B}$ with zero matrices, we can reshape it into an $m$-tuple. Consequently, the trace norm of the correlation matrix can be reformulated as follows (see Appendix G for proof)
$\|\mathcal{C}\|_{\mathrm{tr}}=\max_{\mathcal{C}\in\mathcal{O}(\mathcal{C})}\mathrm{Tr}[\mathcal{C}]\;.$ (27)
The von Neumann’s trace theorem [65] tells us that the maximum is reached when $\boldsymbol{A}^{\prime}=P\boldsymbol{A}$ and $\boldsymbol{B}^{\prime}=Q\boldsymbol{B}$, where $P,Q$ achieve a singular value decomposition (SVD) of $\mathcal{C}$, i.e. $\mathcal{C}=P\Sigma Q^{\mathrm{T}}$. Then, we have
$\|\mathcal{C}\|_{\mathrm{tr}}=\max_{\mathcal{C}\in\mathcal{O}(\mathcal{C})}\mathrm{Tr}[\mathcal{C}]=\sum_{\mu=1}^{r}\braket{A^{\prime}_{\mu}\otimes B^{\prime}_{\mu}}\;.$ (28)
@@
Here, $r=\operatorname{rank}(\mathcal{C})$. Eq. 28 shows that there exist the optimal observables $A^{\prime}_{\mu}$ and $B^{\prime}_{\mu}$ to witness entanglement. Thus, one may immediately get an inequality of correlation function similar to CHSH inequality Eq. 4 for separable state
@@
This implies that, based on the extended Bloch space, entanglement witnesses can be easily generated for generic observables.
7 Concluding remarks
@@ Acknowledgements
This work was supported in part by the National Natural Science Foundation of China(NSFC) under the Grants 12475087, 12235008, the Fundamental Research Funds for the Central Universities and by the University of Chinese Academy of Sciences, and China Postdoctoral Science Foundation funded project No. 2024M753174.
Appendix A The proof of Theorem 1
Given $m$-tuple $\boldsymbol{A}$ and $n$-tuple $\boldsymbol{B}$ with $\sum_{\mu}t_{\mu}^{A}\vec{a}_{\mu}=\sum_{\mu}t_{\mu}^{B}\vec{b}_{\mu}=0$, the correlation matrices of the separable states can be expressed as
@@ Appendix A The proof of Theorem 1
Here, $\vec{\alpha}_{i}$ and $\vec{\beta}_{i}$ are the measurement-induced Bloch vectors (MIBVs) associated with $\boldsymbol{A}$ and $\boldsymbol{B}$ as defined in the main text. Employing the convexity property of trace norm $\|X\|_{\operatorname{tr}}=\mathrm{Tr}\left[\sqrt{X^{\dagger}X}\right]$, we have
@@
Here, we make use of $\sum_{\mu}t_{\mu}^{A}\vec{a}_{\mu}=\sum_{\mu}t_{\mu}^{B}\vec{b}_{\mu}=0$ and ||$\cdot$|| ⋅ | denotes Euclidean norm of a real vector. So, the correlation matrices of the separable states are limited by the size of the measurement-induced Bloch space (MIBS). And we have defined the parameter $\beta_{A}$
$\beta_{A}=\sup_{\rho}\sum_{k=1}^{m}\braket{\bar{A}_{k}}_{\rho}^{2}=\sup_{\vec{\alpha}\in\boldsymbol{B}(A)}|\vec{\alpha}|^{2}$ (34)
similarly for $\beta_{B}$, which yields
$\|\mathcal{C}\|_{\operatorname{tr}}\leq\sqrt{(\vec{t}\,^{A})^{2}/d_{A}^{2}+\beta_{A}}\sqrt{(\vec{t}\,^{A})^{2}/d_{A}^{2}+\beta_{B}}\;.$ (35)
Next, we show that how to obtain an upper bound of $\beta_{A/B}$. Given an $m$-tuple $\boldsymbol{A}$, quantum states can be transformed into the measurement-induced Bloch vector (MIBVs) which constitute the MIBS. It is proved that MIBS is a proper subset of the following hyper-ellipsoid
$\mathcal{B}(\boldsymbol{A})\subseteq\left\{\vec{\alpha}\Big{|}\sqrt{\frac{d}{d\mathrm{Tr}[\rho^{2}]-1}}\vec{\alpha}^{\mathrm{T}}\Omega^{-}\sqrt{\frac{d}{d\mathrm{Tr}[\rho^{2}]-1}}\vec{\alpha}\leq 1\right\}\;.$ (36)
Here, $\Omega^{-}$ denotes the Moore-Penrose inverse of $\Omega$ which is the unique matrix satisfying the following Moore-Penrose conditons [66]
$\Omega\Omega^{-}\Omega=\Omega\;,\;\Omega^{-}\Omega\Omega^{-}=\Omega^{-}\;,$ (37)
$(\Omega\Omega^{-})^{\dagger}=\Omega\Omega^{-}\;,\;(\Omega^{-}\Omega)^{\dagger}=\Omega^{-}\Omega\;.$ (38)
The Moore-Penrose inverse $\Omega^{-}$ coincides with the matrix inverse when $\Omega$ is invertible. Matrix $\Omega$ completely depicts the properties of the ellipsoid defined by Eq. 36, such as the lengths of the semi-axes are given by $\sqrt{\frac{(d\mathrm{Tr}[\rho^{2}]-1)\lambda_{i}}{d}}$, with $\lambda_{i}$ being the eigenvalues of $\Omega$; the eigenvectors of $\Omega$ determines the directions of the semi-axes [67]. Obviously, Bloch space is a special case of $\boldsymbol{A}=\vec{\pi}$. So, for arbitrary $\boldsymbol{A}$, we have
$\beta_{A}\leq\frac{d-1}{d}\lambda_{\max}(\Omega)\;,$ (39)
where $\lambda_{\max}(X)$ is the maximal eigenvalue of matrix $X$. In particular, we have $\beta_{A}=\frac{d-1}{d}\lambda_{\max}(\Omega)$, when Eq. 36 is a hypersphere.
Appendix B The proof of Observation 1